OpenStudy (anonymous):

Find the roots of the polynomial equation. x2 + 2x – 48 = 0 A. –6, 8 B. –4, 12 C. –12, –4 D. –8, 6

OpenStudy (anonymous):

@Mateaus

OpenStudy (anonymous):

First factor then set them equal to 0

OpenStudy (er.mohd.amir):

x2+(8-6)x-48 open bracket factor it

OpenStudy (anonymous):

O_O no

OpenStudy (anonymous):

\[x^2 + 2x -48 = 0\]

OpenStudy (anonymous):

im not good at math:(

OpenStudy (anonymous):

I'm actually really bad at explaining factorials x_x

OpenStudy (anonymous):

(x+8)(x-6)

OpenStudy (anonymous):

then you set them to =0 x+8 = 0 x-6 = 0

OpenStudy (anonymous):

solve for both x's

OpenStudy (anonymous):

help:(..

geerky42 (geerky42):

I am bad at explaining too, but here's my shot;\[(x+a)(x+b) = x^2+(a+b)x+ab,~~~\text{Right?}\] Since you are given \(x^2+2x-48\), you have you have \(a+b=2\) and \(ab=-48\) Can you figure out what \(a\) and \(b\) is?